# Gradient Interference-Aware Graph Coloring for Multitask Learning

### v1

@@ 4.4 Schedule Generation and Execution
Each training step activates exactly one color class; over one period every task in that class receives a gradient update, while conflicting tasks (edges in $E_{\tau}^{(r)}$) are guaranteed not to co‑occur.
4.4.1 Minimum coverage constraint
If the greedy coloring yields a singleton class for a rarely updated task, we replicate that task into each subsequent slot until its update frequency reaches a defined minimum $f_{\min}$.
@@ 4.4.2 Warm-up and annealing
4.5 Overall Algorithm
[…]
4.6 Relationship to Existing Techniques
4.6.1 Versus PCGrad and CAGrad
We avoid projecting or re-weighting gradients. Instead, we temporally separate high-conflict tasks via disjoint scheduling. This eliminates destructive interference without altering step directions.
4.6.2 Versus Selective-Group Updates
Instead of manually defining or tuning task subsets, our scheduler automatically partitions tasks via greedy graph coloring on the conflict graph. This guarantees at most $\Delta+1$ groups and delivers a $\tau$-dependent convergence bound (Section 5.2), all without extra heuristics or manual scheduling.
4.6.3 Versus Re-weighting Techniques
Methods like GradNorm and AdaTask compute losses and back‑propagate gradients for all $K$ tasks at every step, so both memory usage and computational cost grow with $K$. In contrast, our scheduler activates only one color class per step, so memory and compute requirements scale with the size of the active group
@@ 5 Theoretical Analysis
We provide a high-level overview of the theory supporting our scheduler-based approach. For each section, we include expanded proofs and theoretical analysis in the appendix.
5.1 Descent Preservation under $\tau$-Compatibility
Proposition 1.
For any $\tau$-compatible task set $S$
$\Bigl\|\sum_{k\in S}g_{k}\Bigr\|^{2}\;\geq\;(1-\tau)\sum_{k\in S}\|g_{k}\|^{2}.$ (12)
Proof sketch.
Write
$\|{\textstyle\sum_{k}}g_{k}\|^{2}=\sum_{k}\|g_{k}\|^{2}+2\sum_{i<j}\langle g_{i},g_{j}\rangle$ (13)
Because $S$ is $\tau$‑compatible, each inner product satisfies $\langle g_{i},g_{j}\rangle\geq-\tau\|g_{i}\|\|g_{j}\|$. Substitute and bound the cross‑terms with
$-\tau\sum_{i<j}\|g_{i}\|\|g_{j}\|\leq-\tfrac{\tau}{2}\!\sum_{i\neq j}\|g_{i}\|\|g_{j}\|$ (14)
A second application of Cauchy–Schwarz cauchy1821cours converts the mixed sum into
$\tfrac{\tau}{2}(\sum_{k}\|g_{k}\|)^{2}\leq\tau\sum_{k}\|g_{k}\|^{2}.$ (15)
Rearranging yields Equation 13 Yu2020PCGrad . ∎
The inequality introduced in Propostion 1 shows that any task group chosen by our $\tau$-aware scheduler still produces a genuine descent direction for the shared network. Essentially, separating highly conflicting tasks never turns a training step into an ascent step, so the scheduler can avoid interference without negatively impacting progress. The full proof and a clearer explanation is made available in Appendix A.
5.2 Convergence Rate with $\tau$-Dependent Constant
Theorem 1.
Assume each $\mathcal{L}_{k}$ is $L$-smooth, stochastic gradients are unbiased with variance $\leq\sigma^{2}$, and the step size is fixed to $\eta=c/\sqrt{T}$, $c\leq 1/L$. Then the scheduler satisfies
$\min_{1\leq t\leq T}\,\mathbb{E}\!\left[\|\nabla F(\theta_{t})\|^{2}\right]\;\leq\;\frac{2(F_{0}-F^{\star})}{c\sqrt{T}}\,(1+\tau)\;+\;\frac{cL\sigma^{2}}{\sqrt{T}}.$ (16)
Proof sketch.
Start from the descent lemma for smooth non-convex objectives (Ghadimi & Lan, 2013) GhadimiLan2013 . For each step replace $\|g_{t}\|^{2}$ by the lower bound from Proposition 1 with $S=S_{t}$. Summing over $T$ steps and telescoping the function values gives
$\sum_{t=1}^{T}\eta(1-\tfrac{\tau}{2})\,\mathbb{E}\!\bigl[\|\nabla F(\theta_{t})\|^{2}\bigr]\;\leq\;F_{0}-F^{\star}+\frac{\eta^{2}L\sigma^{2}T}{2}.$ (17)
Divide by $T\eta(1-\frac{\tau}{2})$ and optimize in $\eta$ Nesterov2004 to obtain Equation 16. The rate is $O((1+\tau)/\sqrt{T})$; when $\tau=0$ we match the classical bound.
Theorem 1 confirms that the scheduler reaches the classical $O(1/\sqrt{T})$ speed of vanilla SGD while only facing a small $(1+\tau)$ constant for conflict avoidance. Thus our scheduling strategy preserves the well‑known convergence rate even as it enforces low‑interference updates. The full proof is provided in Appendix B.
5.3 Bounded Task Update Staleness
Greedy coloring partitions the task into at most $m\leq\Delta+1$ color groups. We execute those colors cyclically, so a given color (and the task it contains) reappears at most m1m-1 intervening steps. Thus every task’s parameters are never more than m1m-1 iterations old, meeting the usual bounded‑staleness assumption Recht2011Hogwild . Thus the bounded‑staleness bound m1m-1 provides the formal guarantee that our interference‑aware scheduler never lets a task’s parameters drift for more than $\Delta$ iterations, while still ensuring Goal (that highly conflicting tasks are never updated together). Extended proof and practical explanation are available in Appendix C.
5.4 Coloring Period Bound ($\Delta$ + 1)
Because every vertex has at most $\Delta$ neighbors, when you process the vertices one by one the worst case is that all $\Delta$ of those neighbors are already colored and each is using a different color. With the palette $\{1,...,\Delta+1\}$ that still leaves at least one unused color to assign. Repeating this for every vertex ensures the algorithm never needs more than $\Delta+1$ colors in total West2000 ; Diestel2017 . This color bound guarantees that the cycle length of the scheduler is never longer than a property of the conflict graph itself, keeping both memory usage and task-refresh latency predictability small even when the task set grows large. A formal proof and extended explanation is given in Appendix D.
5.5 Baseline Non‑Convex SGD Rate
The smoothness condition lets each SGD step decrease the expected objective up to a controllable error, while the unbiased‑variance term grows only linearly with the number of steps. Balancing these two effects with a step size $\eta=\frac{c}{\sqrt{T}}$ makes the cumulative descent dominate the cumulative noise, yielding the $O\bigl(1/\sqrt{T}\bigr)$ RobbinsMonro1951 ; Nemirovski2009 bound on the smallest expected gradient norm over $T$ iterations. Thus, when $\tau=0$ (no conflicts) the scheduler reduces to ordinary multi-task SGD and recovers the same guarantee Standley2020TaskGrouping ; Ruder2017Overview ; zhang2018overview . This confirms that our methods will never slow training in the absence of interference. A full proof and in-depth explanation is provided in Appendix E.
@@ 5.6 Exact Recovery of the Population Conflict Graph and Task Partition
We show that, after observing gradients for only a modest number of steps, the scheduler can exactly reconstruct the true conflict relations among tasks by average recent gradients (EMA), computing pairwise cosines, thresholding at $\tau$, and coloring the resulting graph. Under natural assumptions about tasks being meaningfully different (separation), noise, and slowly changing gradients (drift), the conflict graph we estimate from finite data agrees, with high probability, with the ideal ’population’ conflict graph. Essentially, the graph we build from averaged gradients matches the true conflict pattern you’d get with unlimited data, meaning the grouping the scheduler uses can effectively recover the ground-truth task partition. The result explains why the scheduler’s group structure is trustworthy and ties the required probe budget (i.e., how many gradient samples must be averaged before building the graph) to interpretable quantities (noise level, margin, number of tasks). Theory supporting this can be found in Appendix F.
5.7 Descent Bounds for Scheduled versus Aggregated Updates
We compare two ways to use the same gradient information from a refresh: (1) a scheduled sequence of per-group steps (i.e., the approach we propose with the scheduler), and (2) a single aggregated step that combines all groups at once. We do this to understand when the scheduler actually helps and by how much. We find that, when different groups’ gradients pull in opposing directions (so adding them together would cancel progress) the scheduler has an advantage. In that case, taking the updates one group at a time is provably better. Our theory guarantees a larger drop in the objective during that refresh than the one-shot step, even though both use the same step size and the same gradients. Theory supporting this can be found in Appendix H.
@@ 6.1 Datasets
Table 1: Information on the datasets utilized in experimentation. (*Samples were removed during preprocessing)
Dataset Main Tasks (+) Aux. Tasks (-) Aux Tasks Modalities Samples
AV-MNIST Digit Classification Digit Parity Audio, Image 56.0k
$M$-IMDb Genre Classification Release Decade Title-Initial Classification Image, Text 25.9k
We evaluate the proposed scheduler alongside numerous baselines and state-of-the-art models across multiple datasets to reliably assess its general performance relative to other approaches. In total, it is evaluated across 6 datasets (see Table 1).
Across all datasets, we incorporate positive and/or negative auxiliary tasks into training. Positive auxiliary tasks share structure or predictive signals with the main tasks (e.g., common features or correlated outputs) and so can improve the learned representations by providing relevant supervision. In contrast, negative auxiliary tasks are uncorrelated or directly conflicting with the main objectives, inducing gradient interference that can slow or degrade primary performance. Including these tasks allows us to stress-test the scheduler’s ability (relative to other models) to suppress harmful tasks and prioritize helpful ones.
6.1.1 NYUv2
The NYU Depth Dataset v2 (NYUv2) Silberman2012 consists of RGB-D indoor scenes with 1,449 densely labeled pairs of RGB and depth images. To demonstrate auxiliary task value in data-scarce conditions, we employ a subset of 250 training samples randomly selected from the original training set.
All tasks utilize RGB images as the sole input modality, with depth maps and surface normals serving as prediction targets rather than input features. A ResNet-18 he2016residual backbone trained from scratch processes the RGB input, with task-specific decoder heads for segmentation (with 32 $\times$ upsampling), depth regression, surface normal regression, and color temperature estimation.
6.1.2 CIFAR-10
The CIFAR-10 krizhevsky2009learning dataset contains 60,000 $32\times 32$ color images across 10 generic classes. To evaluate our interference-aware scheduler in a data-scarce environment where auxiliary tasks provide maximum benefit, we employ a subset of 2,500 training samples (250 per class) from the original 50,000 training images.
For the multi-task learning setup, we set image classification as the main task and construct three auxiliary tasks synthetically from the RGB images. The positive auxiliary tasks include: (1) quadrant localization, where the model predicts which quadrant contains the primary object, and (2) texture classification using Gabor filter responses clustered into 8 texture categories via k-means clustering. The negative auxiliary tasks consist of: (3) corruption-type prediction, where images are artificially corrupted using 15 different corruption types from the ImageNet-C corruption suite hendrycks2019robustness , and (4) rotation angle prediction, where images are rotated by 0°, 90°, 180°, or 270° and the model predicts the rotation angle.
All tasks share a ResNet-18 he2016residual backbone trained from scratch without pretraining, with task-specific heads for each auxiliary task.
6.1.3 AV-MNIST
The AV‑MNIST benchmark vielzeuf2019centralnet pairs MNIST images lecun1998gradient with a log‑mel spectrogram of the corresponding spoken digit from TIDIGITS leonard1993tidigits . It is a synthetic benchmark that has significant noise applied to audio and feature reduction applied to images, making it far more difficult than the original MNIST.
We use all paired samples in our experiments. Our primary task is 10‑way digit classification. Following (Vielzeuf et al., 2018), we encode images with a small 4‑layer convolutional network and spectrograms with a 2‑layer CNN, both built and trained from scratch. These embeddings are projected and fused for processing by a simple MLP in intermediate fusion boulahia2021early ; guarrasi2025systematic , as are the models trained on MM-IMDb and STOCKS. We include only one positive auxiliary class, Digital Parity. This task aims to identify the digits as either even or odd, which has been shown to be a positive auxiliary task for improving representations on MNIST-like datasets Tacchetti2018 ; Mohammadi2020OddOneOut .
6.1.4 MM-IMDb
The MM‑IMDb dataset arevalo2017gatedmultimodalunitsinformation contains 25,959 movies with genre annotations over 23 categories. We extract poster images and plot summaries for every movie in the dataset.
The images and summaries are encoded by a frozen VGG16 simonyan2014very and Google word2vec mikolov2013efficientestimationwordrepresentations model, respectively. Our main task is movie genre prediction. We add one positive auxiliary task, Release Decade, and one negative auxiliary task, the classification of the title’s first word as either a vowel or consonant.
6.1.5 STOCKS
The STOCKS datasets we use, introduced in (Liang et al., 2021) liang2021multibench , contain stock market timeseries data across two categories. Specifically: (1) STOCKS-F&B, which has 14 input and 4 output stocks in the GICS Restaurants or Packaged Food & Meats category MSCI2024GICs , and (2) STOCKS-HEALTH, which contains 56 input and 7 output stocks in the Health Care category.
Every input stock consists of 500 trading days, with the goal of predicting returns over the next day. We discretize the continuous return variable $R$ into three non-overlapping categories: (1) $Low$, where $0\leq R<0.1$, (2) $Medium$, where $0.1\leq R<0.5$, and (3) $High$, where $R\geq 0.5$. Mean Absolute Error (MAE) is calculated by mapping the three classes to numbers ($Low\rightarrow 0$, $Medium\rightarrow 1$, $High\rightarrow 2$) and then deriving MAE as usual. Each input series is encoded by the same CNN-BiLSTM network. This consists of 3 CNNs and 1 BiLSTM Cui2018 .
@@ 6.2 Baseline and State-of-the-Art Comparisons
We evaluate our interference-aware scheduler against 10 different multi-task learning methods across all datasets. These methods include 3 oler baseline models, more recent state-of-the-art gradient-based optimization techniques, and novel combinations of our scheduler with existing methods to demonstrate complementary benefits.
6.2.1 Baseline Models
@@
    Gradnorm Chen2017GradNorm . Balances task learning rates by normalizing gradient magnitudes relative to target loss ratios. This maintains consistent training dynamics across tasks.
@@
    MGDA Sener2018 . Formulates multi-task learning as a multi-objective optimization problem, finding Pareto-optimal solutions lockwood2008pareto ; pareto2014manual through gradient descent in the convex hull of gradients FliegeSvaiter2000 ; Miettinen1999 .
6.2.2 State-of-the-Art Models
  1. 1.
    PCGrad Yu2020PCGrad . Projects conflicting gradients onto orthogonal subspaces when negative cosine similarity is detected, eliminating destructive interference between task gradients.
  2. 2.
    CAGrad Liu2021CAGrad . Extends PCGrad by adaptively adjusting gradient magnitudes based on conflict severity. This proves more nuanced modifications to gradients than binary projection.
  3. 3.
    Adatask Yang2023AdaTask . Dynamically reweighs task losses using relative loss changes, adapting to varying task learning rates during training.
  4. 4.
    FAMO Liu2023FAMO . Fast Adaptive Multitask Optimization dynamically adjusts task weights to equalize each task’s rate of loss improvement. It uses an online, per-step rule (no pairwise gradient ops), adding negligible overhead while remaining robust to loss-scale differences.
  5. 5.
    Fair Resource Allocation in MTL (FairGrad) Ban2024FairMTL . Views the shared update as a limited resource and chooses it to maximize an $\alpha$-fair utility of per-task improvements. The parameter $\alpha$ controls the trade-off between average performance and fairness.
  6. 6.
    Nash-MTL Navon2022NashMTL . Frames multitask training as a bargaining game and computes a scale-invariant weighted combination of task gradients given by the Nash bargaining solution. Weights are obtained by solving a small inner problem (e.g., via CCP) using the gradient Gram matrix. Updates are balanced across tasks.
6.2.3 Scheduler Extension Models
  1. 1.
    Scheduler AdaTask. Combines our interference-aware task selection with AdaTask’s dynamic loss weighting, applying adaptive weights only to scheduler-selected tasks.
  2. 2.
    Scheduler GradNorm Warm Start. Initializes training with GradNorm for stable gradient magnitudes, then transitions to our scheduler after 3 epochs.
  3. 3.
    Scheduler PCGrad. Applied PCGrad’s gradient projection specifically to tasks selected by our scheduler, providing fine-grained conflict resolution within $\tau$-compatible groups.
@@ 6.3.1 Static One-Shot Coloring
We run the greedy graph coloring once at the start of training, freeze the resulting task groups, and never recompute the conflict graph. All other hyperparameters ($\tau$, history length $H$, and update interval $R$) match the full scheduler. As training progresses we expect the fixed coloring to grow stale, mixing tasks whose inference relationships have changed. This ablation isolates the benefit of dynamic recoloring, showing how much performance depends on adapting the schedule to evolving gradient conflicts.

### v2

@@ 4.4 Schedule Generation and Execution
Each training step activates exactly one color class; over one period every task in that class receives a gradient update, while conflicting tasks (edges in $E_{\tau}^{(r)}$) are guaranteed not to co-occur.
4.4.1 Minimum update frequency
If the greedy coloring yields a singleton class for a rarely updated task, we increase its update frequency by duplicating it only into steps whose active color has no conflict edge to that task.
@@ 4.4.2 Warm-up and Annealing
4.5 Time Complexity
The proposed scheduler has a time complexity of $\Theta(K^{2}d)$ per refresh. However, unlike many MTL approaches, our scheduler concentrates its extra work in occasional refreshes. This time complexity therefore becomes $\Theta\!\big(K^{2}d/R\big)$ amortized per training step where $R$ s the refresh period (the number of training steps between conflict-graph rebuilds). It adds non-trivial overhead which grows quadratically with $K$ (number of tasks) but shrinks as $R$ grows. We provide a full analysis of the time complexity in Appendix D and discuss approaches to reducing time complexity under certain conditions in Appendix D.5.
@@ 5 Theoretical Analysis
We discuss some of the main guarantees behind SON-GOKU. For a very brief overview: (1) Updating groups of tasks whose gradients are mostly low-conflict (no internal edges) reduces the objective on average and still achieves the usual $1/\sqrt{T}$ convergence rate; (2) Over a refresh window, scheduling several group updates can beat one mixed update that uses all tasks at once; and (3) With a small number of recent gradient measurements per task (via EMA) and a margin separating conflicts, the estimated conflict graph matches the ideal one, giving a short schedule where every task is updated at least once every $\Delta+1$ steps ($\Delta$ is the maximum number of conflicts for any task). We provide expanded assumptions, definitions, proofs, reasoning, analysis, etc. in Appendix 5.4I.
5.1 Descent Preservation Within a Low-conflict Group
If the active set $S_{t}$ at step $t$ is $\tau$-compatible, then the combined update is a descent direction with a quantitative lower bound:
$\Bigl\|\sum_{k\in S_{t}}g_{k,t}\Bigr\|^{2}\;\geq\;\Bigl(1-\tau\,(|S_{t}|-1)\Bigr)\sum_{k\in S_{t}}\|g_{k,t}\|^{2}$ (12)
Thus the step cannot flip to ascent whenever $\tau(|S_{t}|-1)<1$. This is proved by expanding the polarization identity and controlling cross terms under the $\tau$-compatibility condition (see Appendix E). Essentially, this means that SON-GOKU’s per-step updates are safe when groups are low conflict. The aggregate direction keeps pointing downhill and the cancellation is quantitatively limited by $\tau$ and group size.
5.2 Nonconvex Convergence at the Standard Rate up to a Small Factor
Under standard smoothness and noise conditions (see Appendix I) and with steps $\eta=c/\sqrt{T}$, SON-GOKU achieves the usual nonconvex SGD rate, with a mild $(1+\tau)$ factor that reflects within-group conflict:
$\min_{t<T}\mathbb{E}\,\|\nabla F(\theta_{t})\|^{2}\;\leq\;\frac{2(F_{0}-F^{\star})}{c\sqrt{T}}\,(1+\tau)\;+\;\frac{cL\sigma^{2}}{\sqrt{T}}$ (13)
When $\tau=0$, the constant matches the classical bound (Bottou et al., 2018; Ghadimi & Lan, 2013); as $\tau\to 1$, it at most doubles, matching the intuition that conflict can cancel up to half of the progress. This demonstrates that scheduling does not degrade asymptotic progress. SON-GOKU preserves the $1/\sqrt{T}$ decay of the gradient norm while controlling the constant through the compatibility threshold $\tau$. In other words, we keep the standard rate of SGD and trade a small constant for reduced interference.
5.3 When Scheduled Groups Outperform a Single Mixed Update
We compare two ways to use the same gradients gathered at a refresh: a scheduled sequence of per-group steps (i.e., the scheduler used in SON-GOKU) versus a single aggregated step. Using a telescoping $L$-smooth bound and evaluating both trajectories at a common linearization—i.e., expanding $F$ at the refresh start $\theta_{t_{r}}$ and applying the same first-order model with the same step size—the scheduled bound is never worse and is strictly better when cross-group interaction terms are sufficiently negative (so mixed updates would cancel progress).
Essentially, when different groups’ gradients pull in opposing directions (so adding them together would cancel progress) the scheduler has an advantage. In that case, taking the updates one group at a time is provably better. Our theory guarantees a larger drop in the objective during that refresh than the one-shot step, even though both use the same step size and the same gradients. Under the PL condition, the scheduled path maintains the usual contraction factor and gains a nonnegative extra decrease term over the window.
@@ 5.4 Exact Recovery of the Population Conflict Graph and Task Partition
We show that, after observing gradients for only a modest number of steps, the scheduler can exactly reconstruct the true conflict relations among tasks by averaging recent gradients (EMA), computing pairwise cosines, thresholding at $\tau$, and coloring the resulting graph. Under a separation margin $\gamma$ around the threshold (tasks are meaningfully different), bounded noise, and bounded drift within each refresh window, the conflict graph estimated from finite data agrees, with high probability, with the ideal population conflict graph $G^{\star}\tau$ (defined from the pairwise cosines of the true mean gradients $\{\mu_{i}\}_{i=1}^{K}$ at the start of the refresh window). Equivalently, when the uniform cosine estimation error is below $\gamma$, we have $\widehat{G}\tau=G^{\star}\tau$ and the resulting grouping recovers the ground-truth task partition. This explains why the scheduler’s group structure is trustworthy and ties the required number of recent gradient measurements per task to interpretable quantities such as noise level, margin, and the number of tasks. For example, an effective sample size of $n_{\mathrm{eff}}\;\gtrsim\;\frac{\sigma^{2}}{m_{0}^{2}\,\gamma^{2}}\,\log\!\bigl(K/\delta\bigr)$ suffices in our analysis.
5.5 Scheduling Properties with Few Groups and Bounded Staleness
Welsh-Powell greedy coloring uses at most $\Delta+1$ colors on a graph whose maximum degree is $\Delta$ (Bonamy et al., 2018). Running the colors in a fixed cycle means each task is updated at least once every $m\leq\Delta+1$ steps. Equivalently, no task waits more than $\Delta$ steps between updates (bounded staleness).
This means that the schedule length is controlled by the worst conflict degree $\Delta$ rather than by the total number of tasks $K$. This results in two important benefits: (1) a minimum update-frequency guarantee, since every task receives an update at least once per cycle of length $\leq\Delta+1$; and (2) compatibility with standard bounded-delay conditions used in analyses of asynchronous SGD (e.g., Niu et al. 2011; Lian et al. 2015), with delay parameter at most $\Delta$. When $\Delta\ll K$, we achieve both low interference (few conflicts per step) and low staleness (short update gaps).
@@ 6.1 Datasets
We evaluate across six benchmarks spanning vision, multimodal, and time-series. For each dataset we specify a small set of primary tasks and add positive and negative auxiliaries to stress interference. Architectures are standard backbones (e.g., ResNet-18 for image tasks, CNN/BiLSTM for time-series) with task-specific heads. Full dataset and task definitions, auxiliary construction, and architecture details (including preprocessing and head designs) appear in Appendix J and Table 3.
@@ 6.2 Baseline and State-of-the-Art Comparisons
We compare against loss-weighting (Uniform, GradNorm, AdaTask), multi-objective (MGDA, Nash-MTL, FairGrad), projection/surgery (PCGrad, CAGrad), and fast adaptive weighting (FAMO). We provide short method notes in Appendix K and discuss these approaches in Section 2.
6.3 Scheduler Extension Models
In addition to standalone models, we also evaluate combinations of the scheduler with existing approaches.
@@ 6.3 Scheduler Extension Models
    SON-GOKU ++ AdaTask. Combines our interference-aware task selection with AdaTask’s dynamic loss weighting, applying adaptive weights only to scheduler-selected tasks.
@@
    SON-GOKU ++ GradNorm Warm Start. Initializes training with GradNorm for stable gradient magnitudes, then transitions to our scheduler after 3 epochs.
@@
    SON-GOKU ++ PCGrad. Applied PCGrad’s gradient projection specifically to tasks selected by our scheduler, providing fine-grained conflict resolution within $\tau$-compatible groups.
@@ 6.4.1 Static One-Shot Coloring
We run the greedy graph coloring once at the start of training, freeze the resulting task groups, and never recompute the conflict graph. All other hyperparameters ($\tau$, history length $H$, and update interval $R$) match the full scheduler. As training progresses we expect the fixed coloring to grow stale, mixing tasks whose interference relationships have changed. This ablation isolates the benefit of dynamic recoloring, showing how much performance depends on adapting the schedule to evolving gradient conflicts.